OpenStudy (hannahjane0300):

What are the solutions of the system? Y = X + 5 Y = X^2 - X + 2 A.) (3,14) and (-1, 0) B.) (3, 8) and (-1, 4) C.) (3, 8) and (-1, 0) D.) (-3, 2) and (1, 6)

OpenStudy (dayakar):

y= x+5 -------(1) y = x^2 -x +2------(2) in (1) and (2) left hand sides are equal then right hand sides are equal x^2 -x +2 = x+5

OpenStudy (hannahjane0300):

I'm sorry, I'm still confused.

OpenStudy (dayakar):

do we have two equations is it clear

OpenStudy (hannahjane0300):

Yes.

OpenStudy (dayakar):



OpenStudy (dayakar):

circled are equal then underlined are equal

OpenStudy (hannahjane0300):

Ok.

OpenStudy (dayakar):

x^2 -x +2 = x+5 [ add both sides -x-5]

OpenStudy (dayakar):



OpenStudy (dayakar):

x^2+1x-3x-3 =0 [splitting middle term]

OpenStudy (dayakar):

r u following

OpenStudy (hannahjane0300):

Yes.

OpenStudy (dayakar):

x(x+1)-3(x+1)=0[ taking common] (x+1)(x-3)=0 therefore x+1= 0 or x-3 =0 now find x values

OpenStudy (hannahjane0300):

1 and -3?

OpenStudy (dayakar):

NO x= -1 or x= 3